OpenStudy (anonymous):

what do i do when there is a negative in the front of the sqrt example -sqrt72

OpenStudy (anonymous):

You are aware of complex numbers???

OpenStudy (anonymous):

What topic are you doing??

OpenStudy (phi):

generally people assume the positive square root (the principal value). But with the minus sign, they want you to use the negative value of the square root.

OpenStudy (anonymous):

so how do i do that and im doing a review packet

OpenStudy (phi):

You don't do anything. however, given \[ - \sqrt{72} \] I would try to "simplify" it by factoring 72 into primes: 72= 2*36= 2*3*12=2*3*3*4= 2*3*3*2*2 then look for pairs of numbers: 3*3 and 2*2 to "pull out" of the square root. so \[ -\sqrt{72}= -\sqrt{3*3*2*2*2} =- 3*2\sqrt{2}= -6\sqrt{2} \]

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

what if the problem has a plus or minus sign in front of it

OpenStudy (anonymous):

\[\pm \sqrt{54}\]

OpenStudy (phi):

just leave the plus/minus in the final answer. 54= 2*27= 2*3*9= 2*3*3*3 only 1 pair of numbers: 3*3 so \[\pm 3\sqrt{2*3}= \pm 3\sqrt{6} \]